# Density Fluctuations in the Intracluster Medium: An Attempt to Constrain Viscosity with Cosmological Simulations

### v1

@@ 3 Results
In this section we present the results obtained from the three different clusters and the three different amounts of viscosity employed. We focus first on the intrinsic effect of viscosity with respect to the non-viscous case and then we will try to constrain the viscosity from X-ray observations.
@@ 3.1 Morphology
Viscosity acts by transforming kinetic energy into internal energy, leading to a suppression of the growth of hydrodynamical instabilities and to an increase of the gas temperature (Roediger et al., 2013; Marin-Gilabert et al., 2022). This is expected to produce a strong effect in the morphology of galaxy clusters at small scales (few kpc), where the turbulent cascade is truncated due to the effect of viscosity. Fig. 1 shows the surface density for each of the three clusters and the three different configurations. The inner dotted circle corresponds to $R_{\mathrm{2500c}}$ and the outer dashed circle corresponds to $R_{\mathrm{200c}}$ [^4] In this paper, $R_{\mathrm{2500c}}$ ($R_{\mathrm{200c}}$) is defined as the radius enclosing the region of the cluster with a mean density 2500 (200) times larger than the critical density of the universe in comoving units..
By comparing each cluster independently, we can observe that there are no major changes at large scales (few hundred of kpc - Mpc), where the big structures remain in the same position in the three different cases and the overall gas distribution is very similar. However, the ICM is more homogeneous in the non-viscous case, whereas in the viscous cases we find sharper density discontinuities and more regions with very diffuse gas.
Although at large scale the picture is similar, we find many differences at smaller scales. The lack of mixing due to viscosity leads to a larger amount of small structures that have not been disrupted due to the growth of instabilities. The survival of the clump will depend on different parameters such as the size of the clump, its velocity or the overdensity with respect to the medium (Klein et al., 1994; Pittard et al., 2005; Valentini & Brighenti, 2015), which is beyond of the scope of this paper. The growth or suppression of instabilities will also depend on the amount of viscosity, which depends on the temperature, generating a very complex system difficult to analyse in cosmological simulations. However, its effect can be seen in the amount of clumps observed in Fig. 1.
@@
Because of the strong dependence of viscosity with temperature one might expect that the above-mentioned differences might correlate well with the temperature of the cluster. In Fig. 2 we show a temperature colormap of the nine different simulations. The cases with more viscosity lead to a higher temperature of the diffuse gas, however, the temperature of the dense clumps are very similar in all three cases. Overall, the virial temperature at $z=0$ is around 5% to 10% higher in the viscous cases compared to the ideal ones due to heat dissipation (see table 1). We find again a more homogeneous temperature in the ideal runs, while more fluctuations are present in the viscosity runs. Due to the dependence of viscosity with temperature, viscosity starts to become important at low redshift, when the virial temperature becomes larger than $\sim 10^{7}$ K. In this respect, the merger history of the cluster becomes very important, as mergers behave differently whether we have or we do not have viscosity. This is analysed in full depth in a follow up work. The hotter medium leads to a larger viscosity, which heats up the medium, producing a larger viscosity, entering a runaway cycle which is prevented by the viscous saturation, avoiding unphysical temperatures (see section 2.2).
Figure 1: Density projections for all of our simulations listed in 1. From left to right: MHD only, MHD with $1/3$ of Spitzer viscosity and full Spitzer viscosity at redshift zero. From top to bottom: the runs g55, g16 and g63.
Figure 2: Temperature projections for all of our simulations listed in 1. From left to right: MHD only, MHD with $1/3$ of Spitzer viscosity and full Spitzer viscosity at redshift zero. From top to bottom: the runs g55, g16 and g63.
@@ 3.2.1 Gas distribution
To remove the high-density regions we follow the method introduced in Zhuravleva et al. (2012): we divide the galaxy cluster in spherical shells and compute the density PDF for each of the shells. We calculate the median value of each shell and, alongside it, a threshold value to separate the bulk gas from the high-density gas. A particle is considered to belong to the high-density gas if the following criterion is met:
@@
where $\{n\}$ is the median value of the density in the shell, $f_{\mathrm{cut}}$ is a parameter tuned to select more or less gas as high-density and $\sigma_{10}$ is the standard deviation ($\log_{10}$ based) of the density distribution. $\sigma_{10}$ can be expressed as
@@
Figure 3: Density probability density function (PDF) of a thin shell of 5 kpc of the cluster g55 to show the effect of viscosity in the gas density distribution. The solid lines show the data; the vertical dashed lines indicate the position of the density threshold to separate bulk and high-density gas; and the dash-dotted lines the log-normal distribution fit to the bulk density component of the gas.
@@ 3.2.2 Density fluctuations
Fig. 3, however, only shows a thin slice of cluster g55 to illustrate the effect of viscosity on the gas distribution and the method employed to separate the bulk gas from the high-density gas. To understand the overall effect of viscosity in the whole cluster, Fig. 4 shows the variation of the standard deviation ($\sigma$) of the log-normal fit as a function of the radius for the three different clusters. Assuming the bulk gas follows a log-normal distribution, the value of $\sigma$ can be understood as a density fluctuation measurement, since
@@
where, as before, $\rho_{1}$ and $\rho_{2}$ correspond to the 12th and the 88th percentile of the density, respectively. Increasing the amount of viscosity produces broader distributions along the radius, which translates into larger density fluctuations. Overall, with full Spitzer viscosity we get the largest fluctuations. Comparing the case with $1/3\,\eta$ and the ideal case we still see larger fluctuations in the runs with $1/3\,\eta$, although the differences are not large.
@@ 3.2.3 Density vs velocity fluctuations
where $\langle v_{\mathrm{x}}\rangle$, $\langle v_{\mathrm{y}}\rangle$ and $\langle v_{\mathrm{z}}\rangle$ are the mean velocities in each shell. We then calculate the 1D mach number of each particle as $\mathcal{M}_{\mathrm{1D}}=\delta v/(c_{s}\,\sqrt{3})$. We take the region within $R_{\mathrm{200c}}$ of the cluster, similar to the one taken in Zhuravleva et al. (2023). For better statistics, we take the last 40 snapshots from $z\sim 0.4$ to $z=0$ and calculate the mean value of density and velocity fluctuations of each shell in each snapshot using equations 21 and 22 respectively. For each snapshot we take the same radial bins normalised to $R_{\mathrm{200c}}$. Finally, we take the mean value of the density and velocity fluctuations for each shell over the snapshots and plot $\mathcal{M}_{\mathrm{1D}}$ against $\delta\rho/\rho$, as shown in Fig. 5. Since we have taken the average over a period of time, we cannot take into account the dynamical state of the clusters here.
@@
There is a clear linear trend in all clusters, independently of the amount of viscosity, with a low scatter in all cases (the Pearson’s correlation coefficient is close to 1, indicating a good correlation). Viscosity affects the slope of the linear fit, leading to lower slopes [^5] Note that the slope is the inverse $\eta_{\rho}$.. We observe that density and velocity fluctuations increase further from the center (lighter data points). This reflects the state of the ICM in clusters that keep growing by accretion (Zhuravleva et al., 2023). The values of $\mathcal{M}_{\mathrm{1D}}$ are consistent with direct observations of Perseus (Aharonian et al., 2018), where they found $\mathcal{M}_{\mathrm{3D}}=\sqrt{3}\mathcal{M}_{\mathrm{1D}}\simeq 0.3-0.45$ within the central 100kpc of the cluster ($\sim 0.03R_{\mathrm{200c}}$) and indirect measurements from observations (Lovisari et al., 2024; Dupourqué et al., 2024), with $\mathcal{M}_{\mathrm{3D}}\simeq 0.36-0.41$ for relaxed clusters in the inner regions of the cluster (within $R_{\mathrm{500c}}\sim 0.7R_{\mathrm{200c}}$).
To study in more detail the ratio of density to velocity fluctuations, Fig. 6 shows the dependence of this ratio ($\eta_{\rho}$) as a function of the radius.
Figure 6: Ratio between density fluctuations and velocity fluctuations as a function of the radius within $R_{\mathrm{200c}}$ (same data as in Fig. 5). The solid lines indicate the data from the simulations, while the dashed lines indicate the mean value in each case within $R_{\mathrm{200c}}$. The mean value and the standard deviation in each run is indicated in each panel. From left to right: clusters g55, g16 and g63 for each one of the configurations.
@@ 3.3 Comparison with observations
We have discussed the effects that viscosity has in galaxy clusters, focusing so far on the comparison between viscid and inviscid simulations. However, for a comparison with observations we need to calculate the 3D amplitude fluctuations as a function of the wavenumber $k=1/l$ (where $l$ is the length scale of the fluctuations). To do so, we first need to calculate the density fluctuations of our data elements and then compute the power spectrum. In particle based simulations the computation of the Fourier Transform to calculate the power spectrum is not possible, so we need to interpolate the particle properties into a grid. By doing so, we also avoid spurious effects caused by the voids after removing the high-density regions (see section 3.2.1). In grid simulations, the voids are replaced by the median value of the shell (e.g. Zhuravleva et al., 2023). In our case, the voids are automatically replaced by an average value over the particles of the region due to the interpolation method employed. This is done using the code vortex-p (Vallés-Pérez et al., 2021b, a, Vallés-Pérez et al, subm.), which creates an ad-hoc AMR mesh hierarchy from the density and velocity fields with the ultimate goal of providing a multi-resolution Helmholtz-Hodge and Reynolds decomposition. In this work, we use it exclusively to assign the density and velocity fields onto an AMR mesh, which is done using the same kernel configuration as for evolving the simulation. Once the particle data have been interpolated into the mesh, we calculate the density fluctuations. This is done by decomposing the density field in unperturbed and fluctuating components (Churazov et al., 2012):
@@
are the fluctuations. $\rho_{i}$ is the density of each cell and $\langle\rho\rangle$ is the value of the density profile at that distance of the center (see density profiles in appendix B). In our case, we evaluate the density profile at a given radius to estimate the unperturbed $\rho_{0}(x,y,z)$. For the power spectrum we assume an isotropic fluctuation field $k=\sqrt{k_{x}^{2}+k_{y}^{2}+k_{z}^{2}}$. The 3D amplitude fluctuations can be computed from the power spectrum as
@@
To constrain the amount of viscosity in the ICM, Zhuravleva et al. (2019) suggested a criterion based on studying the slope of density fluctuations. In the cases with higher viscosity the turbulent cascade is stopped earlier, leading to a steeper power spectrum that translates into a steeper 3D density amplitude fluctuations. This was probed by comparing Coma observations with direct numerical simulations (DNS) of hydrodynamic turbulence. They found a level of suppression of at least a factor of 10 with respect to the Spitzer value, depending on the Prandtl number [^6] The Prandtl number is defined as the ratio of momentum to thermal diffusivity (i.e. the ratio between viscosity and thermal conduction)..
@@
Assuming that the motions are subsonic and driven at large scales, one can assume that the fluctuations in the ICM are passively advected by the velocity field (Kunz et al., 2022). Therefore, a comparison with hydrodynamic turbulence simulation is somewhat reasonable. However, the DNS simulations employed in Zhuravleva et al. (2019) use a rather idealised setup, where compressible fluids (important mainly for unrelaxed clusters) or magnetic fields are not considered. They also assumed an isothermal fluid of $T_{e}\sim 8$ keV ($\sim 9.28\times 10^{7}$ K) for Coma. And, since viscosity is highly dependent on the temperature (see equation 9), that means that they used a constant value of viscosity. This might produce the sharp viscous cutoff observed in their results. However, in cosmological simulations, we observe a range of temperatures from $3.3\times 10^{6}$ K to $2.5\times 10^{8}$ K, and therefore, a big range of values for viscosity. This range of values for viscosity at different scales prevent the power spectrum from a sharp cutoff, as observed in Fig. 7 for density and Fig. 15 for velocities power spectrum. As a consequence, the slope of the 3D amplitude density fluctuations is very similar independently of the amount of viscosity. This is analysed in more detail in section 3.4.
@@
It is important to note the maximum observed at $k\sim 2.5R_{\mathrm{2500c}}^{-1}\sim 175$ kpc in the viscous run of cluster g16 in Fig. 7. This feature is not found either in observations or the other simulations, where the amplitude of density fluctuations decrease monotonically with $k$. The reason is the value of $f_{\mathrm{cut}}$ used in equation 18, where we used a value of $f_{\mathrm{cut}}=2.5$, the minimum value suggested in Zhuravleva et al. (2012). This value was reasonable for our cluster g55, where we have a clear separation of bulk gas following a log-normal distribution and a high-density tail (see Fig. 3). However, there is not a clear difference between the bulk gas and high-density regions in the density PDF of cluster g16 (see appendix A), looking more like a log-normal plus power law distribution. Due to its broad PDF, the median value is large and a value of $f_{\mathrm{cut}}=2.5$ sets a threshold in a large density value, producing that very few or no particles are considered as high-density (see Fig. 11). The lack of particles removed due to the choice of $f_{\mathrm{cut}}$ increases the amplitude density fluctuations at the scale of the high-density clumps ($50$ kpc - $200$ kpc). A more reasonable value for this cluster would be $f_{\mathrm{cut}}=0.5$, where the split between bulk and high-density gas is more accurate, as shown in the top panel of Fig. 12. With a lower value of $f_{\mathrm{cut}}$, the maximum observed in the middle panel of Fig. 7 disappears, as can be seen in Fig. 13. However, the same value should be applied to all the analysis for consistency, and that would imply that in cluster g55 a lot of bulk gas would be considered as high-density gas (see bottom panel of Fig. 12). For this reason, we perform our analysis with a value of $f_{\mathrm{cut}}=2.5$, acknowledging that it has a small effect in the viscous run of cluster g16, leading to a factor of $\sim 2$ larger amplitude at the scales of the high-density clumps ($\sim 175$ kpc). However, the slope of the density fluctuations is not affected by the choice of $f_{\mathrm{cut}}$ (see Fig. 13), indicating that the method to constrain viscosity is still robust regardless of the value of $f_{\mathrm{cut}}$.
@@
The code vortex-p also allows us to interpolate the velocity fluctuations field calculated from the $v_{rms}$ as shown in equation 22 and the soundspeed of each particle. Then we calculate the power spectrum of the velocity fluctuations in an analogous way as shown in equation 25
@@
Since viscosity reduces the velocity of the particles (Marin-Gilabert et al., 2022) and, at the same time, increases the temperature (see section 3.1) and, therefore, the soundspeed, one could expect that the $\mathcal{M}_{\mathrm{1D},k}$ decreases with viscosity. However, Fig. 8 shows very similar results among the simulations or even larger amplitude fluctuations in the viscous cases. This apparently counter-intuitive result might be due to the lack of mixing in the viscosity runs. This produces that the most massive structures in the cluster keep their mass for a longer time compered to the non-viscous runs, experiencing a larger gravitational pull and accelerating as they fall into the gravitational well of the cluster. This can be observed also in the kinetic energy spectrum in appendix C, where at large scales the runs with viscosity lead to larger values of the energy spectrum.
Considering density and velocity fluctuations, we can study the relation 12. In Fig. 9 we show the value of $\eta_{\rho,k}$ calculated as $\eta_{\rho,k}=(\delta\rho/\rho)_{k}/\mathcal{M}_{\mathrm{1D},k}$ for different scales. The runs with viscosity tend to have larger values of $\eta_{\rho,k}$. This means that, even though both density and velocity fluctuations are higher in the viscosity runs, they do not compensate for each other and the ratio between the two is still larger than the non-viscous runs. The cluster g55 has values larger than 1, however this is in agreement with previous works for unrelaxed clusters (Gaspari et al., 2014; Simonte et al., 2022). The cluster g16 shows a relation very close to 1, as expected theoretically (Zhuravleva et al., 2014, 2023), although in the case with full viscosity we still see the maximum mentioned previously. In cluster g63 we observe a value close, but lower than 1, also consistent with previous works for a very relaxed cluster.

### v2

@@ 3 Results
In this section we present the results obtained from the three different clusters and the three different values of viscosity employed. We focus first on the intrinsic effect of viscosity with respect to the non-viscous case and then we will try to constrain the viscosity from X-ray observations.
@@ 3.1 Morphology
Viscosity acts by transforming kinetic energy into internal energy, leading to a suppression of the growth of hydrodynamical instabilities and to an increase of the gas temperature (Roediger et al., 2013; Marin-Gilabert et al., 2022). This is expected to produce a strong effect in the morphology of galaxy clusters at small scales (few kpc), where the turbulent cascade is truncated due to the effect of viscosity. Fig. 1 shows the surface density for each of the three clusters and the three different configurations. The inner dotted circle corresponds to $R_{\mathrm{2500c}}$ and the outer dashed circle corresponds to $R_{\mathrm{200c}}$ [^3] In this paper, $R_{\mathrm{2500c}}$ ($R_{\mathrm{200c}}$) is defined as the radius enclosing the region of the cluster with a mean density 2500 (200) times larger than the critical density of the universe..
By comparing each cluster independently, we can observe that there are no major changes at large scales (few hundred kpc - Mpc), where the big structures remain in the same position in the three different cases and the overall gas distribution is very similar. However, the ICM is more homogeneous in the non-viscous case, whereas in the viscous cases we find sharper density discontinuities and more regions with very diffuse gas.
Although at large scale the picture is similar, we find many differences at smaller scales. The lack of mixing due to viscosity leads to more small structures that have not been disrupted due to the growth of instabilities. The survival of the clump will depend on different parameters such as the size of the clump, its velocity or the overdensity with respect to the medium (Klein et al., 1994; Pittard et al., 2005; Valentini & Brighenti, 2015), which is beyond of the scope of this paper. The growth or suppression of instabilities will also depend on the amount of viscosity, which depends on the temperature, generating a very complex system difficult to analyse in cosmological simulations. However, its effect can be seen in the amount of clumps observed in Fig. 1.
@@
Because of the strong dependence of viscosity with temperature one might expect that the above-mentioned differences might correlate well with the temperature of the cluster. In Fig. 2 we show temperature colormaps of the nine different simulations. The cases with more viscosity lead to higher temperatures in the diffuse gas, however, the temperatures of the dense clumps are very similar in all three cases. Overall, the virial temperature at $z=0$ is around 5% to 10% higher in the viscous cases compared to the ideal ones due to heat dissipation (see table 1). We find again a more homogeneous temperature in the ideal runs, while more fluctuations are present in the viscous runs. Due to the dependence of viscosity on temperature, viscosity starts to become important at low redshift, when the virial temperature becomes larger than $\sim 10^{7}$ K. In this respect, the merger history of the cluster becomes very important, as mergers behave differently whether we have or we do not have viscosity. This is analysed in full depth in a follow up work. The hotter medium leads to a larger viscosity, which heats up the medium, producing a larger viscosity, entering a runaway cycle which is prevented by the viscous saturation, avoiding unphysical temperatures (see section 2.2).
Figure 1: Density projections for all of our simulations listed in table 1. From left to right: MHD only, MHD with $1/3$ of Spitzer viscosity and full Spitzer viscosity at redshift zero. From top to bottom: the runs g55, g16 and g63.
Figure 2: Temperature projections for all of our simulations listed in table 1. From left to right: MHD only, MHD with $1/3$ of Spitzer viscosity and full Spitzer viscosity at redshift zero. From top to bottom: the runs g55, g16 and g63.
@@ 3.2.1 Gas distribution
To remove the high-density regions we follow the method introduced in Zhuravleva et al. (2012): we divide the galaxy cluster into spherical shells and compute the density PDF for each of the shells. We calculate the median value of each shell and, alongside it, a threshold value to separate the bulk gas from the high-density gas. A particle is considered to belong to the high-density gas if the following criterion is met:
@@
where $\{n\}$ is the median value of the density in the shell, $f_{\mathrm{cut}}$ is a parameter tuned to select more or less gas as high-density and $\sigma_{10}$ is the standard deviation ($\log_{10}$ based) assuming a log-normal distribution of the density PDF. $\sigma_{10}$ can be expressed as
@@
Figure 3: Density probability density function (PDF) of a thin shell of 5 kpc of the cluster g55 to show the effect of viscosity on the gas density distribution. The solid lines show the data; the vertical dashed lines indicate the values of the density thresholds separating bulk and high-density gas; and the dash-dotted lines the log-normal distribution fit to the bulk density component of the gas.
@@ 3.2.2 Density fluctuations
Fig. 3, however, only shows a thin slice of cluster g55 to illustrate the effect of viscosity on the gas distribution and the method employed to separate the bulk gas from the high-density gas. To understand the overall effect of viscosity in the whole cluster, Fig. 4 shows the variation of the standard deviation ($\sigma$) of the log-normal fit as a function of the radius for the three different clusters. Assuming the bulk gas follows a log-normal distribution, the value of $\sigma$ can be understood as a density fluctuation measurement (i.e. the width of the distribution), since
@@
where, as before, $\rho_{1}$ and $\rho_{2}$ correspond to the 12th and the 88th percentile of the density, respectively. Increasing the amount of viscosity produces broader distributions for all the shells along the radius, which translates into larger density fluctuations. Overall, with full Spitzer viscosity we get the largest fluctuations. Comparing the case with $1/3\,\eta$ and the ideal case we still see larger fluctuations in the runs with $1/3\,\eta$, although the differences are not large.
@@ 3.2.3 Density vs velocity fluctuations
where $\langle v_{\mathrm{x}}\rangle$, $\langle v_{\mathrm{y}}\rangle$ and $\langle v_{\mathrm{z}}\rangle$ are the mean velocities in each shell. We then calculate the 1D mach number of each particle as $\mathcal{M}_{\mathrm{1D}}=\delta v/(c_{s}\,\sqrt{3})$. We take the region within $R_{\mathrm{200c}}$ of the cluster, similar to the one taken in Zhuravleva et al. (2023). For better statistics, we take the last 40 snapshots from $z\sim 0.4$ to $z=0$ and calculate the mean value of density and velocity fluctuations of each shell in each snapshot using equations 17 and 18 respectively. For each snapshot we take the same radial bins normalised to $R_{\mathrm{200c}}$. Finally, we take the mean value of the density and velocity fluctuations for each shell over the snapshots and plot $\mathcal{M}_{\mathrm{1D}}$ against $\delta\rho/\rho$, as shown in Fig. 5. Since we have taken the average over a period of time, we cannot take into account the dynamical state of the clusters here.
@@
There is a clear linear trend in all clusters, independently of the amount of viscosity, with a low scatter in all cases (the Pearson’s correlation coefficient is close to 1, indicating a good correlation). Viscosity reduces the slopes of the linear fits [^4] Note that the slope is the inverse $\eta_{\rho}$.. We observe that density and velocity fluctuations increase further from the center (lighter data points). This reflects the state of the ICM in clusters that keep growing by accretion (Zhuravleva et al., 2023). The values of $\mathcal{M}_{\mathrm{1D}}$ are consistent with direct observations of Perseus (Aharonian et al., 2018), where they found $\mathcal{M}_{\mathrm{3D}}=\sqrt{3}\mathcal{M}_{\mathrm{1D}}\simeq 0.3-0.45$ within the central 100kpc of the cluster ($\sim 0.03R_{\mathrm{200c}}$) and indirect measurements from observations (Lovisari et al., 2024; Dupourqué et al., 2024), with $\mathcal{M}_{\mathrm{3D}}\simeq 0.36-0.41$ for relaxed clusters in the inner regions of the cluster (within $R_{\mathrm{500c}}\sim 0.7R_{\mathrm{200c}}$).
To study in more detail the ratio of density to velocity fluctuations, Fig. 6 shows the radial dependence of this ratio ($\eta_{\rho}$).
Figure 6: Ratio between density fluctuations and velocity fluctuations as a function of the radius within $R_{\mathrm{200c}}$ (same data as in Fig. 5). The solid lines indicate the data from the simulations, while the dashed lines indicate the mean value in each case within $R_{\mathrm{200c}}$. The mean value and the standard deviation calculated over all the shells in each run is indicated in each panel. From left to right: clusters g55, g16 and g63 for each one of the configurations.
@@ 3.3 Comparison with observations
We have discussed the effects that viscosity has in galaxy clusters, focusing so far on the comparison between viscid and inviscid simulations. However, for a comparison with observations we need to calculate the 3D amplitude fluctuations as a function of the wavenumber $k=1/l$ (where $l$ is the length scale of the fluctuations). To do so, we first need to calculate the density fluctuations of our data elements and then compute the power spectrum. In particle based simulations the computation of the Fourier Transform to calculate the power spectrum is not possible, so we need to interpolate the particle properties onto a grid. By doing so, we also avoid spurious effects caused by the voids after removing the high-density regions (see section 3.2.1). In grid simulations, the voids are replaced by the median value of the shell (e.g. Zhuravleva et al., 2023). In our case, the voids are automatically replaced by an average value over the particles of the region due to the interpolation method employed. This is done using the code vortex-p (Vallés-Pérez et al., 2021b, a, 2024), which creates an ad-hoc AMR mesh hierarchy from the density and velocity fields with the ultimate goal of providing a multi-resolution Helmholtz-Hodge and Reynolds decomposition. In this work, we use it exclusively to assign the density and velocity fields onto an AMR mesh, which is done using the same kernel configuration as for evolving the simulation. Once the particle data have been interpolated into the mesh, we calculate the density fluctuations. This is done by decomposing the density field into unperturbed and fluctuating components (Churazov et al., 2012):
@@
are the fluctuations. $\rho_{i}$ is the density of each cell and $\langle\rho\rangle$ is the value of the density profile at that distance from the center (see density profiles in appendix B). In our case, we evaluate the density profile at a given radius to estimate the unperturbed $\rho_{0}(x,y,z)$. For the power spectrum we assume an isotropic fluctuation field that is a function of $k=\sqrt{k_{x}^{2}+k_{y}^{2}+k_{z}^{2}}$. The 3D amplitude fluctuations can be computed from the power spectrum as
@@
To constrain the amount of viscosity in the ICM, Zhuravleva et al. (2019) suggested a criterion based on studying the slope of density fluctuations spectrum. In the cases with higher viscosity the turbulent cascade is stopped earlier, leading to a steeper power spectrum that translates into steeper 3D density amplitude fluctuations. This was probed by comparing Coma observations with direct numerical simulations (DNS) of hydrodynamic turbulence. They found a level of suppression of at least a factor of 10 with respect to the Spitzer value, depending on the Prandtl number [^5] The Prandtl number is defined as the ratio of momentum to thermal diffusivity (i.e. the ratio between viscosity and thermal conduction)..
@@
Assuming that the motions are subsonic and driven at large scales, one can assume that the fluctuations in the ICM are passively advected by the velocity field (Kunz et al., 2022). Therefore, a comparison with hydrodynamic turbulence simulation is somewhat reasonable. However, the DNS simulations employed in Zhuravleva et al. (2019) use a rather idealised setup, where compressible fluids (important mainly for unrelaxed clusters) or magnetic fields are not considered. They also assumed an isothermal fluid of $T_{e}\sim 8$ keV ($\sim 9.28\times 10^{7}$ K) for Coma. And, since viscosity is highly dependent on the temperature (see equation 9), that means that they used a constant value of viscosity. This might produce the sharp viscous cutoff observed in their results. However, in cosmological simulations, we observe a range of temperatures from $3.3\times 10^{6}$ K to $2.5\times 10^{8}$ K, and therefore, a big range of values for viscosity. This range of values for viscosity at different scales prevents the power spectrum from having a sharp cutoff, as observed in Fig. 7 for density and Fig. 15 for velocities power spectrum. As a consequence, the slope of the 3D amplitude of density fluctuations is insensitive to the amount of viscosity. This is investigated in more detail in section 3.4.
@@
It is important to note the maximum observed at $k\sim 2.5R_{\mathrm{2500c}}^{-1}\sim 175$ kpc in the viscous runs of clusters g16 and g63 in Fig. 7. This feature is not found either in observations or the other simulations, where the amplitude of density fluctuations decreases monotonically with $k$. The reason is the value of $f_{\mathrm{cut}}$ used in equation 14, where we used a value of $f_{\mathrm{cut}}=2.5$, the minimum value suggested in Zhuravleva et al. (2012). This value was reasonable for our cluster g55, where we have a clear separation of bulk gas following a log-normal distribution and a high-density tail (see Fig. 3). However, there is no clear separation between the bulk gas and high-density regions in the density PDF of cluster g16 (see appendix A) and g63, which looks more like a log-normal plus power law distribution. Due to its broad PDF, the median value is large and a value of $f_{\mathrm{cut}}=2.5$ sets a threshold at a large density value, causing very few or no particles to be considered as high-density (see Fig. 11). The small fraction of particles removed due to the choice of $f_{\mathrm{cut}}$ increases the amplitude density fluctuations at the scale of the high-density clumps ($50$ kpc - $200$ kpc). A more reasonable value for this cluster would be $f_{\mathrm{cut}}=0.5$, where the split between bulk and high-density gas is more appropriate, as shown in the top panel of Fig. 12. With a lower value of $f_{\mathrm{cut}}$, the maximum observed in the middle panel of Fig. 7 disappears, as can be seen in Fig. 13. However, the same value should be applied to all the analysis for consistency, and that would imply that in cluster g55 a lot of bulk gas would be considered as high-density gas (see bottom panel of Fig. 12). For this reason, we perform our analysis with a value of $f_{\mathrm{cut}}=2.5$, acknowledging that it has an impact on the viscous runs of clusters g16 and g63, leading to a factor of $\sim 2$ larger amplitude at the scales of the high-density clumps ($\sim 175$ kpc). However, the slope of the density fluctuations is not affected by the choice of $f_{\mathrm{cut}}$ (see Fig. 13), indicating that the method to constrain viscosity is still robust regardless of the value of $f_{\mathrm{cut}}$.
@@
The code vortex-p also allows us to interpolate the velocity fluctuation field calculated from $v_{rms}$ as shown in equation 18 and the soundspeed of each particle. Then we calculate the power spectrum of the velocity fluctuations in an analogous way as shown in equation 21
@@
Since viscosity reduces the velocity of the particles (Marin-Gilabert et al., 2022) and, at the same time, increases the temperature (see section 3.1) and, therefore, the soundspeed, one could expect that the $\mathcal{M}_{\mathrm{1D},k}$ decreases with viscosity. However, Fig. 8 shows very similar results among the simulations or even larger amplitude fluctuations in the viscous cases. This apparently counter-intuitive result might be due to the lack of mixing in the viscosity runs. This causes that the most massive structures in the cluster keep their mass for a longer time compered to the non-viscous runs, experiencing a larger gravitational pull and accelerating as they fall into the gravitational well of the cluster before being stripped. This can be observed also in the kinetic energy spectrum in appendix C, where at large scales the runs with viscosity lead to larger values of the energy spectrum.
Considering density and velocity fluctuations, we can study the relationship between the density and velocity fluctuations of equation 10. In Fig. 9 we show the value of $\eta_{\rho,k}$ calculated as $\eta_{\rho,k}=(\delta\rho/\rho)_{k}/\mathcal{M}_{\mathrm{1D},k}$ for different scales. The runs with greater viscosity tend to have larger values of $\eta_{\rho,k}$. This means that, even though both density and velocity fluctuations are higher in the viscous runs, they do not compensate for each other and the ratio between the two is still larger than for the non-viscous runs. The cluster g55 has values larger than 1, however this is in agreement with previous work for unrelaxed clusters (Gaspari et al., 2014; Simonte et al., 2022). The cluster g16 shows a relation very close to 1, as expected theoretically (Zhuravleva et al., 2014, 2023), although in the case with full viscosity we still see the maximum analysed in appendix A. In cluster g63 we observe a value close to, but lower than 1, also consistent with previous work for a very relaxed cluster.